20.0g NH3 and 50.0g O2 are allowed to react which is the limiting reagent?
1
Expert's answer
2020-03-19T02:35:28-0400

"4NH_3(g) +3 O_2(g) \longrightarrow 2N_2(g)+6H_2O(g)"

molar mass of NH3= 14+3=17 "\frac{g}{mole}" , given mass ammonia=20 g


"moles of NH_3= \frac{given mass}{molar mass}=\frac{20}{17}=1.18 mole"

molar mass of O2=32 g/mole


"mole of O_2= \frac{given mass}{molar mass}= \frac{50}{32}=1.56 mole"

from the above reaction it is clear that

4 mole of ammonia react with 3 mole of oxygen

so, 1 mole of ammonia will react with "=\frac{3}{4} mole of O_2= 0.75 mole" of O2

so 1.18 mole of NH3 will react with "1.18\times.75=0.885" mole of O2

so from above situation NH3 is limiting reagent